OpenStudy (anonymous):

find the area of the bermuda triangle if the sides of the triangle have the approximately lengths 841 miles, 926 miles, and 1305 miles?

OpenStudy (anonymous):

Do you know not know how to find the area of a triangle with 3 given side lengths? @pparth

OpenStudy (mathstudent55):

Use Heron's formula

OpenStudy (anonymous):

but how do you solve it out! can you help me on this one

OpenStudy (anonymous):

Exactly what @mathstudent55 said. Heron's formula is the key to finding the area of a triangle with 3 given sides. There is other ways too like using a combination of cosine/sine laws but the most straightforward approach here would be to simply use Heron's formula.

OpenStudy (anonymous):

can you help me on this one as a example, so I know how to do the rest?

OpenStudy (anonymous):

Here is what Heron's Formula looks like:\[\bf A = \sqrt{s(s-a)(s-b)(s-c)}\]Where\[\bf s=\frac{ a+b+c }{ 2 }\]and a, b, c are the sides of the triangle.

OpenStudy (anonymous):

Now just plug in the numbers 841, 926, 1305, for any of a, b, c and solve for the Area. @pparth

OpenStudy (anonymous):

Looks like he mysteriously disappeared in the triangle

OpenStudy (anonymous):

hmmmmmmmmmmmmmmmmmmmmmmm

OpenStudy (anonymous):

anyone knows: (5,-4pie/3) its sayings to find a polar coordinates of the point

OpenStudy (anonymous):

Make a new question for this and I'll help you out. @pparth

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

5,-4pie/3) its sayings to find a polar coordinates of the point